# arxiv-appendix

### v1

@@ Abstract
Motivated by privacy concerns in sequential decision-making on sensitive data, we address the challenge of nonparametric contextual multi-armed bandits (MAB) under local differential privacy (LDP). We develop a uniform-confidence-bound-type estimator, showing its minimax optimality supported by a matching minimax lower bound. We further consider the case where auxiliary datasets are available, subject also to (possibly heterogeneous) LDP constraints. Under the widely-used covariate shift framework, we propose a jump-start scheme to effectively utilize the auxiliary data, the minimax optimality of which is further established by a matching lower bound. Comprehensive experiments on both synthetic and real-world datasets validate our theoretical results and underscore the effectiveness of the proposed methods.
Keywords: local differential privacy, contextual multi-armed bandit, covariate shift
@@ 1 Introduction
Contextual multi-armed bandit (MAB) (e.g. Lu et al., 2010; Zhou, 2016) is a versatile and general model with applications in various practical domains, such as personalized recommendations (e.g. Li et al., 2010), clinical trials (e.g. Ameko et al., 2020), and portfolio management (e.g. Cannelli et al., 2023). However, existing algorithms for these applications often rely heavily on sensitive user data. For instance, clinical trial results often include detailed physical and biometric information about patients, raising concerns about potential data leakage when these results are shared with parties beyond specific hospitals. To address these concerns, differential privacy (DP) (Dwork et al., 2006) has emerged as the gold standard for protecting user data. Depending on the availability of a central server that has access to all information, the notion of DP is further categorized into central differential privacy (CDP) and local differential privacy (LDP) (e.g. Kairouz et al., 2014; Duchi et al., 2018). Vast literature has considered private contextual MAB under the CDP framework where a central server can store user data (e.g. Kusner et al., 2015; Shariff and Sheffet, 2018; Dubey and Pentland, 2020; Han et al., 2022; Wang et al., 2022; He et al., 2022; Chakraborty et al., 2024; Chen et al., 2025a).
@@
In some scenarios, such a central server does not exist. For instance, mobile users may distrust companies and avoid sharing sensitive information directly. In such cases, LDP serves as an effective privacy-preserving framework. In fact, compared to CDP, LDP is applicable to a greater number of real-world scenarios and is more widely deployed in industry (Erlingsson et al., 2014; Apple, 2017; Tang et al., 2017; Yang et al., 2024). Contextual MAB has also been extensively studied under LDP settings (e.g. Dubey and Pentland, 2020; Zheng et al., 2020; Han et al., 2021; Chowdhury and Zhou, 2022; Garcelon et al., 2022; Charisopoulos et al., 2023; Huang et al., 2023; Li et al., 2024; Chen et al., 2025b), primarily focusing on reward functions with parametric assumptions, such as linear structures. A similar problem, where the action space is a continuous real space, have also been explored under both parametric (Zhao et al., 2024) and nonparametric assumptions (Chen et al., 2023). However, no prior research has addressed the nonparametric LDP contextual MAB problem.
@@
With the vast amount of data resources in the big data era, we often have access to additional data sources besides information from the target problem. Transfer learning (TL) (e.g. Cai and Wei, 2021; Li et al., 2022; Cai and Pu, 2024; Li et al., 2024), a promising area of research in machine learning and statistics, aims to improve performance in a target domain by leveraging knowledge from related source domains, achieving significant success across various applications. Such improvement usually relies on the assumption that the target and source distributions share certain similarities, such as the regression function (e.g. Cai and Wei, 2021; Pathak et al., 2022) or the sparsity structure (e.g. Li et al., 2022). As demonstrated by Zhang and Bareinboim (2017), Suk and Kpotufe (2021), Cai et al. (2024), among others, such source knowledge is particularly effective in MAB, as it can significantly boost the performance of policies in early stages that would otherwise incur high regret. With additional need of preserving privacy, no existing literature further investigates private contextual MAB with knowledge transfer.
@@
Identifying these gaps, this work considers contextual MAB under LDP and addresses the following three key questions: (i) What is the theoretical limit of LDP contextual MAB without parametric assumption? (ii) Can auxiliary data extend this limit? (iii) Can effective algorithms be designed to solve contextual MAB with LDP while also incorporating auxiliary data?
@@
When addressing these questions, we impose privacy constraints on auxiliary datasets as well as the target one, allowing auxiliary datasets to follow distributions that are different from the target, but adhering to the covariate shift framework. In particular, the framework assumes identical reward functions but allows different marginal distributions. It is particularly suitable when there exists an objectively homogeneous conditional relationship (the reward function) across several parties with population heterogeneity (e.g. Suk and Kpotufe, 2021; Cai et al., 2024). For example, the expected outcomes of a clinical trial represent an objective relationship that remains consistent when conditioned on patient features. However, the distribution of patient features may vary across different cooperating medical institutions.
@@
With the aforementioned setup, our contributions are summarized as follows: (i) We formalize the problem of contextual MAB under LDP and extend it by introducing auxiliary datasets under covariate shift. (ii) We establish minimax lower bounds on the regret, accounting for varying levels of privacy and the extent of covariate shift. (iii) We propose a uniform-confidence-bound-type policy for LDP contextual MAB, along with a jump-start scheme to leverage auxiliary data effectively. (iv) We derive a high-probability upper bound on the regret achieved by our policy, which is shown to be minimax near-optimal. (v) We provide extensive numerical results on both synthetic and real datasets to validate our theoretical findings and demonstrate the practical utility of our methodology.
@@
In Section 2, we introduce the problem of contextual MAB under LDP and present the corresponding methodologies and theoretical results. The content related to auxiliary datasets is covered in Section 3. Numerical results and discussions are provided in Sections 4 and 5, respectively. All proofs and detailed descriptions of the experiments are included in the appendix.
Notation.
@@
For any vector $x$, let $x^{i}$ denote the $i$-th element of $x$. For $1\leq p<\infty$, the $L_{p}$-norm of $x=(x^{1},\ldots,x^{d})^{\top}$ is defined by $\|x\|_{p}:=(|x^{1}|^{p}+\cdots+|x^{d}|^{p})^{1/p}$. We use the notation $a_{n}\lesssim b_{n}$ and $a_{n}\gtrsim b_{n}$ to denote that there exist positive constants $n_{1}\in\mathbb{N}$, $c$ and $c^{\prime}$ such that $a_{n}\leq cb_{n}$ and $a_{n}\geq c^{\prime}b_{n}$, respectively, for all $n\geq n_{1}$. We denote $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. Let $a\vee b=\max(a,b)$ and $a\wedge b=\min(a,b)$. For any set $A\subset\mathbb{R}^{d}$, the diameter of $A$ is defined by $\mathrm{diam}(A):=\sup_{x,x^{\prime}\in A}\|x-x^{\prime}\|_{2}$. Let $f_{1}\circ f_{2}$ represent the composition of functions $f_{1}$ and $f_{2}$. Denote the $k$-composition of function $f$ by $f^{\circ k}$. Let $A\times B$ be the Cartesian product of sets, where $A\in\mathcal{X}_{1}$ and $B\in\mathcal{X}_{2}$ for potentially different domains $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$. For measure $\mathrm{P}$ on $\mathcal{X}_{1}$ and $\mathrm{Q}$ on $\mathcal{X}_{2}$, define the product measure $\mathrm{P}\otimes\mathrm{Q}$ on $\mathcal{X}_{1}\times\mathcal{X}_{2}$ as $\mathrm{P}\otimes\mathrm{Q}(A\times B)=\mathrm{P}(A)\mathrm{Q}(B)$. For a positive integer $k$, denote the $k$-fold product measure on $\mathcal{X}_{1}^{k}$ as $\mathrm{P}^{k}$. Let the standard Laplace random variable have probability density function $e^{-|x|}/2$ for $x\in\mathbb{R}$. Let $\operatorname{Unif}(\mathcal{X})$ be the uniform distribution over any domain $\mathcal{X}$. A ball whose center and radius are $x$ and $r\in(0,+\infty)$, respectively, is denoted as $B(x,r)$. For a positive integer $K$, denote $[K]=\{1,2,\ldots,K\}$ and $[0]=\emptyset$.
2 Locally Private Contextual Bandits
@@ 2.1 Preliminaries
Privacy.
We first rigorously define the notion of LDP.
@@
This formulation is widely adopted (e.g. Duchi et al., 2018), with the final estimator constructed based only on the private data $\tilde{Z}_{1},\ldots,\tilde{Z}_{n}$. The term sequentially interactive refers to the privacy mechanisms having access to the privatized historical data. This is particularly suitable for describing the streaming nature of MAB.
Contextual multi-armed bandits.
Let domain $\mathcal{X}=[0,1]^{d}$, number of arms $K\in\mathbb{Z}_{+}$ and $\mathrm{P}$ be a probability measure supported on $\mathcal{X}\times[0,1]^{K}$, generating $(X^{\mathrm{P}},Y^{\mathrm{P},(1)},\ldots,Y^{\mathrm{P},(K)})$. Denote the time horizon by $[n_{\mathrm{P}}]$. At time $t\in[n_{\mathrm{P}}]$, based on the covariate $X_{t}^{\mathrm{P}}\in\mathcal{X}$ drawn from the marginal distribution $\mathrm{P}_{X}$, an arm $k\in[K]$ is selected to receive a random reward $Y_{t}^{\mathrm{P},(k)}\in[0,1]$ associated with the chosen $k$, whose value is drawn according to the conditional distribution $\mathrm{P}_{Y^{P,(k)}|X_{t}^{\mathrm{P}}}$. Given $X_{t}^{\mathrm{P}}$, let the conditional expectation of $Y_{t}^{\mathrm{P},(k)}$ be
@@
Figure 1: Illustration of the learning process.
Let $\pi^{*}$ denote the oracle optimal policy with access to full knowledge of the reward functions $\{f_{k}\}_{k=1}^{K}$, namely $\pi^{*}(x)\in\mathop{\mathrm{argmax}}_{k\in[K]}f_{k}(x)$. Our main objective is to design a policy $\pi$ minimizing the regret defined as
@@
2.2 Minimax Regret Bound
In this section, we investigate the minimax rate of the regret in the problems of contextual MAB subject to LDP. The rate is materialized through a lower bound in Theorem 2.5 and an upper bound Theorem 2.6. The specific class of distributions considered is denoted by $\Lambda(K,\beta)$, i.e.
@@ 2.2 Minimax Regret Bound
Assume that the reward functions $\{f_{k}\}_{k=1}^{K}$ are Lipschitz continuous, i.e. there exists an absolute constant $C_{L}>0$ such that
$\big{|}f_{k}(x)-f_{k}(x^{\prime})\big{|}\leq C_{L}\|x-x^{\prime}\|_{2},\quad\forall x,x^{\prime}\in\mathcal{X},\quad k\in[K].$
@@
Assume that the marginal density $\mathrm{P}_{X}$ is bounded, i.e. there exist absolute constants $\overline{c}>\underline{c}>0$ such that $\underline{c}r^{d}\leq\mathrm{P}_{X}(B(x,r))\leq\overline{c}r^{d}$ for any $x\in\mathcal{X}$ and $r\in(0,1]$.
@@
Assume that the reward functions $\{f_{k}\}_{k=1}^{K}$ satisfy the margin condition, i.e. there exist absolute constants $\beta,C_{\beta}>0$ such that
@@
Assumptions 2.2 and 2.3 are standard in nonparametric statistics literature (e.g. Audibert and Tsybakov, 2007; Samworth, 2012; Chaudhuri and Dasgupta, 2014). 2.4 upper-bounds the probability of the event, where the best arm is hard to distinguish. The larger $\beta$ is, the larger the separation henceforth the easier the problem. This way of characterizing the difficulty of the problem is widely used in the bandit literature (e.g. Perchet and Rigollet, 2013; Suk and Kpotufe, 2021; Cai et al., 2024). As noted by Perchet and Rigollet (2013), when $\beta>d$—that is, when the separation is excessively significant—one of the arms becomes uniformly dominant across $\mathcal{X}$. Then the problem reduces to a static MAB, which is not our focus. Consequently, we only consider $\beta\leq d$.
@@
where $c>0$ is an absolute constant depending only on $d$, $C_{L}$ and $\beta$. In particular, when $0<\varepsilon\leq 1$, it holds with constant $c^{\prime}>0$ that
@@
The proof of Theorem 2.5 can be found in Appendix LABEL:sec:Proof-of-Lower-Bound. In the following, we present a high-probability upper bound on the regret, which can be achieved by Algorithm 2 detailed in Section 2.3. The proof of Theorem 2.6 is provided in Appendix LABEL:sec:proof-upper-bound.
@@
Consider the class of distributions $\Lambda(K,\beta)$ in (2) and the class of LDP policies $\Pi(\varepsilon)$. Suppose $\mathrm{P}\in\Lambda(K,\beta)$. Then, we have $(1)$ the policy $\pi$ given by Algorithm 2 satisfies $\pi\in\Pi(\varepsilon)$ and $(2)$ with probability at least $1-n_{\mathrm{P}}^{-2}$,
@@
where $C>0$ is an absolute constant depending only on $d$, $C_{L}$ and $\beta$. If in addition that $0<\varepsilon\leq 1$, then it holds with constant $C^{\prime}>0$ that
@@
We first compare Theorems 2.5 and 2.6 in a popular case, where the number of arms $K=O(1)$. Up to logarithmic factors, in the challenging, high-privacy regime $\varepsilon\in(0,1]$, Theorems 2.5 and 2.6 together lead to the minimax rate for the regret
@@
The regret (7) is decreasing with respect to both $\varepsilon$ and $\beta$, as larger $\varepsilon$ and $\beta$ correspond to an easier problem. Observing the left-hand side of (7), the two terms correspond to private and non-private rates, where the private rate always dominates with $\varepsilon\in(0,1]$.
We investigate the private and non-private rates separately. The non-private term in (7) is $n_{\mathrm{P}}^{1-\frac{1+\beta}{2+d}}$, consistent with the standard rate for nonparametric contextual MAB under Lipschitz continuity (e.g. Perchet and Rigollet, 2013; Suk and Kpotufe, 2021; Cai et al., 2024). As for the private term, the average regret over $n_{\mathrm{P}}$ target data is $\left(n_{\mathrm{P}}\varepsilon^{2}\right)^{-\frac{1+\beta}{2+2d}}$, aligning with known convergence rates for generalization error of nonparametric classification under LDP constraints (Berrett and Butucea, 2019). Compared to the non-private average regret which is $n_{\mathrm{P}}^{-\frac{1+\beta}{2+d}}$, the LDP rate suffers an extra factor of $d$ in the exponent, thus exhibiting a more severe curse of dimensionality—an effect commonly observed in previous LDP studies (Berrett et al., 2021; Sart, 2023; Györfi and Kroll, 2023).
We conclude this subsection with discussions regarding the gap between the upper and lower bounds in terms of the logarithmic factors, the number of arms $K$ and the privacy budget $\varepsilon$. The additional logarithmic term arises due to the high probability argument we use. As for $K$, the upper bound (5) depends on $K$, while the lower bound (3) does not. Such disagreement between the upper and lower bounds in terms of $K$ is observed in the literature (e.g. Perchet and Rigollet, 2013; Suk and Kpotufe, 2021). A more refined analysis on closing the gap regarding $K$ remains a challenging open problem. We conjecture that this roots in the severe dependence on the dimension of utility loss under privacy constraints. Increasing $K$ amplifies the variance of estimation, and this amplification is further exacerbated by the injected privacy noise. For moderate $\varepsilon$, there remains a gap between $e^{\varepsilon}-1$ dependence in the lower bound (3) and $\varepsilon$ in the upper bound (5). We conjecture that the lower bound is sharp and a different policy is on demand. Such phenomenon is commonly observed in LDP literature (Györfi and Kroll, 2023; Xu et al., 2023; Ma and Yang, 2024), with rates in the moderate $\epsilon$ regime only studied in simple hypothesis testing literature as for now (e.g. Pensia et al., 2023).
2.3 Methodology

### v2

# Locally Private Nonparametric Contextual Multi-armed Bandits with Transfer Learning

@@ Abstract
Motivated by privacy concerns in sequential decision-making on sensitive data, we address the challenging problem of nonparametric contextual multi-armed bandits (MAB) under local differential privacy (LDP). Via a novelly designed LDP-compatible confidence bound, we propose an algorithm that achieves near-optimal regret performance, whose optimality is further supported by a newly derived minimax lower bound. We further consider the case of private transfer learning where auxiliary datasets are available, subject also to (heterogeneous) LDP constraints. Under the widely-used covariate shift framework, we propose a jump-start scheme and a novel reweighted LDP-compatible estimator and confidence bound, which effectively combine and utilize information from heterogeneous auxiliary data. The minimax optimality of the algorithm is further established by a matching lower bound. Comprehensive experiments on both synthetic and real-world datasets validate our theoretical results and underscore the effectiveness of the proposed methods.
Keywords: local differential privacy, contextual multi-armed bandit, transfer learning, covariate shift
@@ 1 Introduction
Contextual multi-armed bandit (MAB) (e.g. Lu et al.,, 2010; Zhou,, 2016) is a versatile and general framework for sequential decision-makings and has been widely deployed in various practical domains, such as personalized recommendations (e.g. Li et al.,, 2010), clinical trials (e.g. Ameko et al.,, 2020), and portfolio management (e.g. Cannelli et al.,, 2023). However, the contextual information in many applications often consists of sensitive user data. For example, clinical trials may include detailed physical and biometric information about patients, while recommendation systems may hold demographics and purchase/view histories information of users. It thus naturally raises privacy concerns given potential data leakage of the sensitive contextual information in MAB.
@@
To address the information security concerns, differential privacy (DP) (Dwork et al.,, 2006) has emerged as the gold standard for protecting user data. Depending on the availability of a central server that has access to all information, the notion of DP can be further categorized into central differential privacy (CDP) and local differential privacy (LDP) (e.g. Kairouz et al.,, 2014; Duchi et al.,, 2018). In the literature, under a parametric assumption on the reward functions, many works have considered private contextual MAB under the CDP setting where a trusted central server can store user data (e.g. Kusner et al.,, 2015; Shariff and Sheffet,, 2018; Dubey and Pentland,, 2020; Wang et al.,, 2022; Chakraborty et al.,, 2024; Chen et al.,, 2025).
@@
However, in many practical scenarios, such a trusted central server may not exist and users may prefer to avoid directly sharing any sensitive information with the server. In such cases, LDP serves as an effective privacy-preserving framework. In fact, compared to CDP, LDP is more widely deployed in the industry due to its greater applicability (Erlingsson et al.,, 2014; Apple,, 2017; Tang et al.,, 2017; Yang et al.,, 2024). In the literature, contextual MAB has also been studied under the LDP setting (e.g. Zheng et al.,, 2020; Han et al.,, 2021; Charisopoulos et al.,, 2023; Huang et al.,, 2023; Li et al.,, 2024; Zhao et al.,, 2024), though existing works also primarily focus on parametric reward functions, such as linear and generalized linear models. Indeed, to our knowledge, no prior work has addressed the problem of nonparametric contextual MAB under LDP constraints.
@@
In the era of big data, the decision makers (referred to as server henceforth), such as financial, pharmaceutical and tech companies, often have access to additional data sources (i.e. auxiliary data) besides information from the target problem. This motivates transfer learning (TL) (e.g. Cai and Wei,, 2021; Li et al.,, 2022; Cai and Pu,, 2024), a promising area of research in machine learning and statistics, which aims to improve performance in a target domain by leveraging knowledge from related source domains. Substantial improvement can be achieved via TL when the target and source problems share certain similarities, such as regression function (e.g. Cai and Wei,, 2021; Pathak et al.,, 2022) or sparsity structure (e.g. Li et al.,, 2022). Importantly, existing works show that TL can effectively leverage auxiliary data and improve regret in both parametric (e.g. Zhang and Bareinboim,, 2017) and nonparametric contextual MAB (e.g. Suk and Kpotufe,, 2021; Cai et al.,, 2024), as it can significantly boost the performance of policies in early stages that would otherwise incur high regret. However, with the additional need of preserving privacy, no existing work has investigated private contextual MAB with knowledge transfer.
@@
Identifying these gaps, our work considers contextual MAB under the LDP constraints and aims to address the following three key questions: (i) What is the fundamental limit of nonparametric contextual MAB under LDP? (ii) Can TL with auxiliary data extend this limit? (iii) Can effective algorithms be designed to solve contextual MAB with LDP while also incorporating auxiliary data?
@@
Our framework allows LDP constraints on both target and auxiliary data. Aligned with the TL literature on contextual MAB (e.g. Suk and Kpotufe,, 2021; Cai et al.,, 2024), we follow the covariate shift framework, where the target and source MAB have the same reward functions but their contextual information may follow different marginal distributions. This setting is suitable when there exists an objectively homogeneous conditional relationship (i.e. the reward function) across several parties with population heterogeneity As a concrete example, the expected outcomes of a clinical trial represent an objective relationship that remains consistent when conditioned on patient features. However, the distribution of patient features may vary across different cooperating medical institutions.
@@
With the aforementioned setup, our contributions are summarized as follows: (i) We formalize the problem of nonparametric contextual MAB under LDP and further extend it to private transfer learning by introducing auxiliary datasets under covariate shift. (ii) We derive minimax lower bounds on the regret, accounting for varying levels of privacy and the extent of covariate shift. (iii) Based on a novelly designed LDP-compatible confidence bound, we propose an efficient policy for LDP contextual MAB, along with a jump-start scheme to further leverage auxiliary data. (iv) We derive a high-probability regret upper bound for the proposed policy, which is near-optimal and matches the minimax lower bound. (v) We conduct extensive numerical experiments on both synthetic and real data to validate our theoretical findings and demonstrate the practical utility of our methodology.
@@
In Section 2, we introduce the problem of nonparametric contextual MAB with LDP and present the proposed methods and theoretical results. We further extend the problem to private TL with auxiliary data in Section 3. Numerical results, including real data applications, and a conclusion with discussions are provided in Sections 4 and 5, respectively. All technical proofs and detailed descriptions of the numerical experiments are included in the supplement.
@@
Notation. For any vector $x$, let $x^{i}$ denote the $i$-th element of $x$. For $1\leq p<\infty$, the $L_{p}$-norm of $x=(x^{1},\ldots,x^{d})^{\top}$ is defined by $\|x\|_{p}:=(|x^{1}|^{p}+\cdots+|x^{d}|^{p})^{1/p}$. We use the notation $a_{n}\lesssim b_{n}$ and $a_{n}\gtrsim b_{n}$ to denote that there exist positive constants $n_{1}\in\mathbb{N}$, $c$ and $c^{\prime}$ such that $a_{n}\leq cb_{n}$ and $a_{n}\geq c^{\prime}b_{n}$, respectively, for all $n\geq n_{1}$. We denote $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. Let $a\vee b=\max(a,b)$ and $a\wedge b=\min(a,b)$. For any set $A\subset\mathbb{R}^{d}$, the diameter of $A$ is defined by $\mathrm{diam}(A):=\sup_{x,x^{\prime}\in A}\|x-x^{\prime}\|_{2}$. Let $f_{1}\circ f_{2}$ represent the composition of functions $f_{1}$ and $f_{2}$. Denote the $k$-composition of function $f$ by $f^{\circ k}$. Let $A\times B$ be the Cartesian product of sets, where $A\in\mathcal{X}_{1}$ and $B\in\mathcal{X}_{2}$ for potentially different domains $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$. For measure $\mathrm{P}$ on $\mathcal{X}_{1}$ and $\mathrm{Q}$ on $\mathcal{X}_{2}$, define the product measure $\mathrm{P}\otimes\mathrm{Q}$ on $\mathcal{X}_{1}\times\mathcal{X}_{2}$ as $\mathrm{P}\otimes\mathrm{Q}(A\times B)=\mathrm{P}(A)\mathrm{Q}(B)$. For a positive integer $k$, denote the $k$-fold product measure on $\mathcal{X}_{1}^{k}$ as $\mathrm{P}^{k}$. Let the standard Laplace random variable have probability density function $e^{-|x|}/2$ for $x\in\mathbb{R}$. Let $\operatorname{Unif}(\mathcal{X})$ be the uniform distribution over any domain $\mathcal{X}$. A ball whose center and radius are $x$ and $r\in(0,+\infty)$, respectively, is denoted as $B(x,r)$. Denote $[K]=\{1,2,\ldots,K\}$ and $[0]=\varnothing$.
2 Locally Private Nonparametric Contextual Bandits
@@ 2.1 Preliminaries
Privacy. We first rigorously define the notion of LDP.
@@
This LDP formulation is widely adopted (e.g. Duchi et al.,, 2018), with the statistical procedure operating based only on the private data $\tilde{Z}_{1},\ldots,\tilde{Z}_{n}$. The term sequentially interactive refers to the privacy mechanisms having access to the privatized historical data, which is particularly suitable for describing the sequential nature of bandit problems.
Contextual multi-armed bandits. Let domain $\mathcal{X}=[0,1]^{d}$, number of arms $K\in\mathbb{Z}_{+}$ and $\mathrm{P}$ be a probability measure supported on $\mathcal{X}\times[0,1]^{K}$, generating $(X^{\mathrm{P}},Y^{\mathrm{P},(1)},\ldots,Y^{\mathrm{P},(K)})$. Denote the time horizon by $[n_{\mathrm{P}}]$. At time $t\in[n_{\mathrm{P}}]$ (i.e. for the $t$-th user), based on the covariate $X_{t}^{\mathrm{P}}\in\mathcal{X}$ drawn from the marginal distribution $\mathrm{P}_{X}$, an arm $k\in[K]$ is selected and one receives a random reward $Y_{t}^{\mathrm{P},(k)}\in[0,1]$ associated with the chosen $k$, whose value is drawn according to the conditional distribution $\mathrm{P}_{Y^{\mathrm{P},(k)}|X_{t}^{\mathrm{P}}}$. Given $X_{t}^{\mathrm{P}}$, let the conditional expectation of $Y_{t}^{\mathrm{P},(k)}$ be
@@
Figure 1: Illustration of the learning process. To achieve LDP, the server only receives privatized information $\tilde{Z}^{\mathrm{P}}_{t}$, while the context $X_{t}^{\mathrm{P}}$, the pulled arm $\pi_{t}(X_{t}^{\mathrm{P}})$, and the reward $Y_{t}^{\mathrm{P}}$ remains at the user end. The same applies to the auxiliary data.
Let $\pi^{*}$ denote the oracle optimal policy with access to full knowledge of the reward functions $\{f_{k}\}_{k=1}^{K}$, namely $\pi^{*}(x)\in\mathop{\mathrm{argmax}}_{k\in[K]}f_{k}(x)$. Our main objective is to design a LDP-preserving policy $\pi\in\Pi(\varepsilon)$ minimizing the regret defined as
@@
2.2 Minimax Optimal Regret Bound
In this section, we investigate the minimax optimal rate of the regret in the problems of contextual MAB subject to LDP. The rate is materialized through a lower bound in Theorem 2.5 and an upper bound in Theorem 2.6. The specific class of distributions considered is denoted by $\Lambda(K,\beta)$, i.e.
@@ 2.2 Minimax Optimal Regret Bound
The reward functions $\{f_{k}\}_{k=1}^{K}$ are Lipschitz continuous, i.e. there exists an absolute constant $C_{L}>0$ such that
$\big{|}f_{k}(x)-f_{k}(x^{\prime})\big{|}\leq C_{L}\|x-x^{\prime}\|_{2},\text{ for all }x,x^{\prime}\in\mathcal{X}\text{ and }k\in[K].$
@@
The marginal density $\mathrm{P}_{X}$ is bounded, i.e. there exist absolute constants $\overline{c}>\underline{c}>0$ such that $\underline{c}r^{d}\leq\mathrm{P}_{X}(B(x,r))\leq\overline{c}r^{d}$ for any $x\in\mathcal{X}$ and $r\in(0,1]$.
@@
The reward functions $\{f_{k}\}_{k=1}^{K}$ satisfy the margin condition, i.e. there exist absolute constants $\beta,C_{\beta}>0$ such that
@@
Assumptions 2.2 and 2.3 are standard in the nonparametric statistics literature (e.g. Audibert and Tsybakov,, 2007; Samworth,, 2012; Chaudhuri and Dasgupta,, 2014). 2.4 upper bounds the probability of the event where the best arm is hard to distinguish. The larger $\beta$ is, the larger the separation and hence the easier the problem. This characterization of the difficulty of the problem is widely used in the bandit literature (e.g. Rigollet and Zeevi,, 2010; Perchet and Rigollet,, 2013; Suk and Kpotufe,, 2021; Cai et al.,, 2024). As noted by Perchet and Rigollet, (2013), when $\beta>d$—that is, when the separation is excessively large—one of the arms becomes uniformly dominant across $\mathcal{X}$. The problem then reduces to a static MAB, which is not our focus. Consequently, we only consider $\beta\leq d$.
@@
where $c>0$ is an absolute constant depending only on $d$, $C_{L}$ and $\beta$. In particular, when $0<\varepsilon\leq 1$, it holds with an absolute constant $c^{\prime}>0$ that
@@
The proof of Theorem 2.5 can be found in Section LABEL:sec:Proof-of-Lower-Bound of the supplement. To accompany the lower bound, in the following, we further present a high-probability upper bound on the regret, which can be achieved by a novel nonparametric LDP bandit algorithm proposed in Section 2.3 (Algorithm 2) later. The proof of Theorem 2.6 is provided in Section LABEL:sec:proof-upper-bound.
@@
Consider the class of distributions $\Lambda(K,\beta)$ in (2) and the class of LDP policies $\Pi(\varepsilon)$. Suppose $\mathrm{P}\in\Lambda(K,\beta)$. Then, we have that the policy $\pi$ given by Algorithm 2 satisfies $\pi\in\Pi(\varepsilon)$ and with probability at least $1-n_{\mathrm{P}}^{-2}$,
@@
where $C>0$ is an absolute constant depending only on $d$, $C_{L}$ and $\beta$. If in addition that $0<\varepsilon\leq 1$, then it holds with an absolute constant $C^{\prime}>0$ that
@@
We first compare Theorems 2.5 and 2.6 for the case widely encountered in practice, where the number of arms $K=O(1)$. Up to logarithmic factors, in the challenging, high-privacy regime $\varepsilon\in(0,1]$, Theorems 2.5 and 2.6 together lead to the minimax rate for the regret
@@
The regret in (7) is a decreasing function of both $\varepsilon$ and $\beta$, which is intuitive as larger $\varepsilon$ and $\beta$ correspond to an easier problem. Observing the left-hand side of (7), the two terms correspond to private and non-private rates, where the private rate always dominates with $\varepsilon\in(0,1]$.
We now provide a detailed discussion on the private and non-private rates in (7). The non-private term is $n_{\mathrm{P}}^{1-\frac{1+\beta}{2+d}}$, consistent with the standard rate for nonparametric contextual MAB under Lipschitz continuity (e.g. Perchet and Rigollet,, 2013; Suk and Kpotufe,, 2021; Cai et al.,, 2024). As for the private term in (7), the average regret over $n_{\mathrm{P}}$ target data is $\left(n_{\mathrm{P}}\varepsilon^{2}\right)^{-\frac{1+\beta}{2+2d}}$, aligning with known convergence rates for generalization error of nonparametric classification under LDP constraints (Berrett and Butucea,, 2019). Compared to the non-private average regret, which is $n_{\mathrm{P}}^{-\frac{1+\beta}{2+d}}$, the LDP rate suffers an extra factor of $d$ in the exponent, thus exhibiting a more severe curse of dimensionality—an effect commonly observed in previous LDP studies (Berrett et al.,, 2021; Sart,, 2023; Györfi and Kroll,, 2023).
We conclude this subsection with discussions regarding the gap between the upper and lower bounds in terms of the logarithmic factors, the number of arms $K$ and the privacy budget $\varepsilon$. The additional logarithmic term arises due to the high probability argument we use. As for $K$, the upper bound (5) depends on $K$, while the lower bound (3) does not. Such disagreement between the upper and lower bounds in terms of $K$ is also observed in the literature for non-private nonparametric MAB (e.g. Perchet and Rigollet,, 2013; Suk and Kpotufe,, 2021). Note that in practice, the number of arms $K$ is typically fixed, which makes this gap less relevant. A more refined analysis on closing the gap regarding $K$ remains a challenging open problem. For moderate $\varepsilon$, there is a gap between $e^{\varepsilon}-1$ dependence in the lower bound (3) and $\varepsilon$ in the upper bound (5). We conjecture that the lower bound is sharp and a different policy is needed to match it. Such phenomenon is commonly observed in the LDP literature (Györfi and Kroll,, 2023; Xu et al.,, 2023; Ma and Yang,, 2024), with rates in the moderate $\varepsilon$ regime only studied in the simple hypothesis testing setting (e.g. Pensia et al.,, 2023).
2.3 Upper Bound Methodology